OpenStudy (anonymous):

@AlexandervonHumboldt2 mind guiding me through this last fill in the blank? Thanks!

OpenStudy (anonymous):

2t + 8 is equal to or greater than -4(t + 1) t is equal to or greater than ___

OpenStudy (alexandervonhumboldt2):

ok so we have 2t+8>=-4*(t+1) distribute: 2t+8>=-4t-4 add 4t to both sides 2t+4t+8>=-4t+4t-4 6t+8-8>=-4-8 6t>=-12 now divide both sides by 6 and flip the sind as one side is negative: 6t/6<=-12/6 t<=-2

OpenStudy (alexandervonhumboldt2):

wait

OpenStudy (alexandervonhumboldt2):

noonoon dont flip the sign

OpenStudy (anonymous):

Well, now you have me majorly confused haha

OpenStudy (alexandervonhumboldt2):

so it would be 6t>=-12 t>=-2

OpenStudy (anonymous):

OH OKAY haha thanks

OpenStudy (alexandervonhumboldt2):

\(\Huge\color{black}{☺☻☺☻}\)